# Amplifying Main Memory-Based Timing Covert and Side Channels using Processing-in-Memory Operations

### v3

@@ Abstract
The adoption of processing-in-memory (PiM) architectures has been gaining momentum because they provide high performance and low energy consumption by alleviating the data movement bottleneck. Yet, the security of such architectures has not been thoroughly explored. The adoption of PiM solutions provides a new way to directly access main memory, which malicious user applications can potentially exploit. We show that this new way to access main memory opens opportunities for high-throughput timing attack vectors that are hard-to-mitigate without significant performance overhead.
We introduce IMPACT, a set of high-throughput main memory-based timing attacks that leverage characteristics of PiM architectures to establish covert and side channels. IMPACT enables high-throughput communication and private information leakage by exploiting the shared DRAM row buffer. To achieve high throughput, IMPACT (i) eliminates expensive cache bypassing steps required by processor-centric main memory and cache-based timing covert and side-channel attacks and (ii) leverages the intrinsic parallelism of PiM operations. We showcase two applications of IMPACT. First, we build two covert-channel attack variants that run on the host CPU and leverage different PiM approaches (i.e., processing-near-memory and processing-using-memory) to gain direct and fast access to main memory and establish high-throughput communication covert channels. Second, we showcase a side-channel attack that leaks private information of concurrently running victim applications that are accelerated with PiM. Our results demonstrate that (i) our covert channels achieve 12.87 Mb/s and 14.16 Mb/s communication throughput, respectively, which is up to 4.91$\times$ and 5.41$\times$ faster than the state-of-the-art main memory-based covert channels, and (ii) our side-channel attack allows the attacker to leak secrets with a low error rate. To avoid such covert and side channels in emerging PiM systems, we propose and evaluate three defenses that fundamentally eliminate the timing channel.
11footnotetext: K. Kanellopoulos and F. N. Bostancı are co-primary authors.
@@ 1 Introduction
Data movement between the main memory and the processor is a critical bottleneck present in today’s computing systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Hence, both industry and academia have invested significant efforts to bring computation closer to where data resides, creating a paradigm called processing-in-memory (PiM)  [13, 23, 8, 24, 25, 26, 11, 27, 19, 28, 1, 2]. PiM solutions provide significant performance improvement and energy efficiency by reducing data movement.
@@
PiM solutions take either of two approaches: (i) processing-near-memory (PnM), where the computation happens in a computation unit placed on the same chip as the memory array [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 21, 39, 40, 41, 42], which also appears in real-world products (e.g., UPMEM [43], AxDIMM [41], AiM [44], HB-PNM [45] and FIMDRAM [46]) and (ii) processing-using-memory (PuM) where the memory array performs the computation by leveraging the analog computation capability of DRAM cells [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. Prior works on PnM and PuM clearly demonstrate that these emerging architectures promise great performance and energy efficiency benefits. In addition, the existence of real products that implement such techniques [43, 41, 44, 45, 46] indicates that these techniques are likely to be used in many real systems in the near future. Therefore, it is important to investigate the security aspect of such techniques and implement countermeasures to avoid potential widespread security vulnerabilities that their deployment can cause. Unfortunately, no prior work analyzes and evaluates the security of emerging PiM architectures against timing covert and side-channel attacks.
@@
In this work, we analyze existing PiM architectures and show that the adoption of PiM solutions creates opportunities for high-throughput main memory-based attacks that are hard-to-mitigate without significant performance overheads. This is possible due to two main reasons: (i) PiM architectures provide direct access to main memory, which is essential for their performance and energy efficiency benefits, and (ii) a large set of existing practical defense mechanisms are not applicable in these PiM architectures.
@@
1. Direct Access to Main Memory. High-throughput main memory-based timing side- and covert-channel attacks need to access main memory with high frequency to observe or change the states of shared main memory structures (e.g., DRAM row buffers). In today’s systems, it is difficult to access main memory with high frequency due to the deep cache hierarchies that filter the memory requests of attacker processes. [^1] We provide analysis as to why prior attack primitives that bypass caches do not provide fast and reliable direct main memory access in §3.2.
@@
The adoption of PiM techniques changes how applications and computation units access main memory: (i) due to area and thermal dissipation limitations, compute units in PnM architectures have small or no caches that allow the user space applications to bypass caches more easily (i.e., with fewer memory accesses), and (ii) PuM architectures provide the ability to offload specialized PiM instructions (e.g., bulk data copy and initialization, and bitwise/arithmetic operations) to the main memory that completely bypass the deep cache hierarchies of modern processors. As a result, PiM architectures provide user space applications with fast and reliable direct main memory access that can be exploited by attackers for high-throughput timing covert- and side-channel attacks.
@@
2. Hard-to-Mitigate with Practical Defenses. A fundamental solution to memory-based timing attacks is to eliminate the timing channel itself. Prior defenses explore fundamental solutions such as memory partitioning and enforcing constant-time memory accesses. However, these defenses incur significantly high performance overheads (as shown in Section 6). Therefore, several prior works [63, 64, 65, 66] explore practical defenses and detection mechanisms that restrict access to timing channels with cache management-based methods, such as restricting the access to cache management instructions (e.g., clflush instruction in x86 systems [63]) and detecting timing attacks based on cache access patterns (e.g., a high number of cache misses can indicate a timing attack) [64, 65, 66]. These mechanisms are inapplicable to PiM-based attacks because these attacks completely bypass the cache hierarchy.
@@
Based on our analysis, we introduce IMPACT, a set of high-throughput In-Memory Processing-based timing Attacks that leverage the direct and fast main memory accesses enabled by PiM architectures. Using the direct access via PiM operations, IMPACT observes and exploits the shared DRAM row buffer. We demonstrate that IMPACT achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing covert- and side-channel attacks and (ii) leveraging the intrinsic parallelism of PiM operations. We showcase this finding on two covert- and one side-channel attacks. To protect future systems against IMPACT covert and side channels, we propose three defense mechanisms that fundamentally eliminate the channel itself.
@@
IMPACT-PnM Covert Channel Variant. We build a PnM-based covert channel that uses PiM-enabled instructions [67] to directly access the main memory. The sender bypasses the locality monitor employed by the system and offloads PiM-enabled instructions for computation directly inside the main memory. The receiver measures the time it takes to access a specific set of rows to retrieve the message encoded with row buffer hits and conflicts. IMPACT-PnM achieves 12.87 Mb/s communication throughput, which is up to 4.91$\times$ faster than the state-of-the-art main memory-based attack [68].
IMPACT-PuM Covert Channel Variant. We build a PuM-based covert channel that leverages the inherent parallelism in main memory using bulk copy and initialization techniques, specifically RowClone [52], to quickly transmit messages through the shared row buffer states. This way, IMPACT-PuM transmits 16-bit large messages 14$\times$ faster than IMPACT-PnM and achieves communication throughput of 14.16 Mb/s (up to 5.41$\times$ faster than the state-of-the-art [68]).
@@
IMPACT Side-Channel Attack. We build a PnM-based IMPACT side-channel attack that exploits PiM operations to leak private information belonging to a concurrently running application based on its memory access characteristics. We showcase this side-channel attack on read mapping, one of the fundamental applications in DNA sequence analysis where preserving the privacy of the human genomic data is crucial [69, 70, 71, 72, 73]. Our side-channel attack leverages PiM operations and leaks the private characteristics of a user’s sample genome by observing the memory access patterns of the read mapping application. Our evaluation demonstrates that our side-channel attack leaks the properties of a sample genome at a throughput of 7.57 Mb/s with 96% accuracy.
@@
Defense Mechanisms. Based on our security analysis, we discuss and evaluate potential defense mechanisms that can mitigate IMPACT attacks by eliminating the timing channel. We discuss the performance versus security trade-offs of these defenses. We demonstrate that the two best-performing defense mechanisms lead to high performance overheads of 26% and 15% on average across five data-intensive workloads.
@@
  • To the best of our knowledge, this is the first work that analyzes and evaluates the security of emerging PiM architectures against timing covert and side-channel attacks. We introduce IMPACT, a new set of high-throughput main memory-based timing attacks that leverage PiM operations to gain fast and reliable access to main memory.
  • We showcase two IMPACT covert channel variants, IMPACT-PnM and IMPACT-PuM that exploit PnM and PuM techniques, respectively. First, IMPACT-PnM allows the sender and the receiver to exchange messages using PiM-enabled instructions. Second, IMPACT-PuM further leverages the parallelism provided by bulk copy PuM techniques to improve the communication throughput. We show that both attacks outperform the state-of-the-art main-memory-based covert-channel attacks in terms of communication throughput.
  • We showcase an IMPACT side-channel attack on read mapping by leveraging PnM instructions. We show that our side-channel attack enables the attacker to leak sensitive information on the victim’s sample genome with high throughput and a low error rate.
  • We evaluate and discuss the trade-offs of three defense mechanisms to protect PiM-enabled systems against IMPACT’s covert- and side-channel attacks.
@@ 2.1 DRAM Organization and Operation
Organization. Fig. 1 shows the hierarchical organization of a modern DRAM-based main memory. The memory controller connects to a DRAM module over a memory channel. A module contains one or more DRAM ranks that time-share the memory channel. A rank consists of multiple DRAM chips that operate in lock-step. Each DRAM chip contains multiple DRAM banks that can be accessed independently. A DRAM bank is organized as a two-dimensional array of DRAM cells, where a row of cells is called a DRAM row. A DRAM cell consists of 1) a storage capacitor, which stores one bit of information in the form of electrical charge, and 2) an access transistor, which connects the capacitor to the row buffer through a bitline controlled by a wordline.
@@
Operation. To perform a read or write operation, the memory controller first needs to open a row, i.e., copy the data of the cells in the row to the row buffer. To open a row, the memory controller issues an activation command to a bank by specifying the row address to open. After activation is complete, the memory controller issues either read and write commands to read or write a DRAM word within the activated row. Subsequent accesses to the same row can be served quickly without fetching the data to the row buffer, creating a row hit [74]. To access data from another DRAM row in the same bank, the memory controller must first close the currently open row by issuing a bank precharge command, creating a row conflict/miss.
@@ 2.2 Processing-in-Memory
Processing-in-memory (PiM) moves computation closer to where the data resides, alleviating the data movement bottleneck between memory and the processor [13, 23, 8, 24, 25, 26, 11, 27, 19, 28, 1, 2]. There are two main PiM approaches: 1) processing-near-memory (PnM) and 2) processing-using-memory (PuM). PnM place compute logic near memory arrays (e.g., processing capability in the memory controller and logic layer(s) of 3D-stacked memory [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 21, 39, 40, 41, 42]). PnM logic can execute application regions (e.g., instructions [67, 32], functions [28], application threads [20, 54]) depending on the design. PnM computation units typically employ small caches due to 1) area constraints and 2) the ineffectiveness of caches in improving the performance of a wide range of memory-intensive applications. Therefore, various systems [32, 67, 28] offload the memory-intensive application regions on the PnM units, while executing more cache-friendly application regions on the host CPU with a deeper cache hierarchy. PuM uses analog operational properties of the memory circuitry (e.g., DRAM) to enable massively parallel in-memory computation [47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62].
@@ 2.3 Main Memory Timing Attacks
Main memory timing attacks exploit several shared structures in the main memory. One such structure is the DRAM row buffer that acts as an internal cache inside DRAM banks. An attacker can understand if a row is recently accessed (i.e., already in the row buffer) or not by measuring the time it takes to access that row. Using this observation, DRAMA [68] builds a side-channel attack to leak keystrokes in a real system. Several prior works [68, 75, 76, 77, 78] leak DRAM address mapping functions. Other main memory-based attacks exploit the memory bus [79], and memory deduplication [80, 81] and (de)compression [82, 83, 84, 85, 86, 87, 88, 89, 90] latencies.
@@ 3 A Case for PiM-based Timing Attacks
The adoption of processing-in-memory (PiM) architectures has been gaining momentum in recent years because they provide high performance and low energy consumption by reducing data movement. Therefore, it is timely and critical to analyze and evaluate the security of PiM architectures to prevent any vulnerabilities from causing widespread susceptibility as PiM architectures become widely available. Unfortunately, no prior work analyzes and evaluates the security of emerging PiM architectures against timing covert and side-channel attacks. After a thorough analysis of existing PiM architectures, we show that the adoption of PiM solutions creates opportunities for high-throughput main memory-based timing attacks for two main reasons. First, the adoption of PiM solutions provides user space applications with fast and reliable direct access to main memory. Second, timing attacks in PiM architectures can be more difficult to detect and mitigate due to commonly employed mechanisms, like cache monitoring mechanisms that track the cache miss ratio during runtime to detect anomalies [64, 65, 66], being either inapplicable or performance-inefficient in PiM architectures.

### v4

# Revisiting Main Memory-Based Covert and Side Channel Attacks in the Context of Processing-in-Memory

@@ Abstract
We introduce IMPACT, a set of high-throughput main memory-based timing attacks that leverage characteristics of processing-in-memory (PiM) architectures to establish covert and side channels. IMPACT enables high-throughput communication and private information leakage by exploiting the shared DRAM row buffer. To achieve high throughput, IMPACT (i) eliminates expensive cache bypassing steps required by processor-centric memory-based timing attacks and (ii) leverages the intrinsic parallelism of PiM operations. We showcase two applications of IMPACT. First, we build two covert channels that leverage different PiM approaches (i.e., processing-near-memory and processing-using-memory) to establish high-throughput covert communication channels. Our covert channels achieve 8.2 Mb/s and 14.8 Mb/s communication throughput, respectively, which is 3.6$\times$ and 6.5$\times$ higher than the state-of-the-art main memory-based covert channel. Second, we showcase a side-channel attack that leaks private information of concurrently-running victim applications with a low error rate. Our source-code is openly and freely available at https://github.com/CMU-SAFARI/IMPACT.
11footnotetext: F. N. Bostancı and K. Kanellopoulos are co-primary authors.
@@ 1 Introduction
Data movement between computation units (e.g., CPUs, GPUs, ASICs) and main memory (e.g., DRAM) is a major performance and energy bottleneck in current processor-centric computing systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], and is expected to worsen due to the increasing data intensiveness of modern applications, e.g., machine learning [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33] and genomics [34, 35, 36, 37, 38, 39, 40, 41, 42]. To mitigate the overheads caused by data movement, various works propose Processing-in-Memory (PiM) architectures [13, 43, 8, 44, 45, 46, 11, 47, 19, 48, 1, 2]. There are two main approaches to PiM [49, 10, 8]: (i) Processing-near-Memory (PnM) [50, 51, 52, 53, 37, 54, 55, 38, 56, 57, 21, 58, 59, 60, 26], where computation logic is added near the memory arrays (e.g., in a DRAM chip, next to each bank, or at the logic layer of a 3D-stacked memory [61, 62, 63, 64, 65, 66, 20, 67, 21, 11, 68, 25, 69, 35, 26]); and (ii) Processing-using-Memory (PuM) [70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85], where computation is performed by exploiting the analog operational properties of the memory circuitry.
@@
Prior works from industry and academia demonstrate the large performance and energy benefits of PiM architectures. A set of PiM techniques are already implemented in real products [86, 60, 87, 88, 89] and many more are expected to be adopted in the near future. Therefore, it is important to investigate the security aspects of such techniques and implement countermeasures to avoid potential widespread security vulnerabilities that their deployment can cause. Unfortunately, no prior work analyzes and evaluates the security of emerging PiM architectures against timing covert- and side-channel attacks.
@@
In this work, we analyze PiM architectures and show that the adoption of PiM architectures creates opportunities for critical main memory-based timing attacks due to two reasons. First, to eliminate data movement, PiM architectures provide direct access to main memory, which is a key building block for high-throughput main memory-based timing attacks. Second, defenses against these attacks either incur high performance overheads or are not applicable to PiM architectures.
@@
1. Direct Access to Main Memory. Main memory-based timing attacks establish covert communication channels or leak secrets by observing and changing the shared main memory states (e.g., DRAM row buffer states). To achieve high throughput, timing attacks require fast and reliable access to main memory. In today’s systems, it is difficult to access main memory with high frequency due to the deep cache hierarchies that filter the memory requests and incur additional latency overheads [90]. [^1] We provide analysis as to why prior attack primitives that bypass caches do not provide fast and reliable enough direct main memory access in §3.2.
@@
The adoption of PiM techniques changes how applications and computation units access main memory: (i) due to area and thermal dissipation limitations, compute units in PnM architectures have small or no caches that allow the userspace applications to bypass caches more easily (i.e., with fewer instructions and lower latency), and (ii) PuM architectures provide the ability to offload to main memory specialized PiM instructions (e.g., bulk data copy and initialization [75, 91, 73, 92, 85], and bitwise/arithmetic operations [76, 93, 73, 79, 78, 58, 85, 81, 80, 82]) that completely bypass the deep cache hierarchies of modern processors. As a result, PiM architectures provide userspace applications with fast and reliable direct main memory access that can be exploited by attackers for high-throughput timing covert- and side-channel attacks.
@@
2. Hard-to-Mitigate with Practical Defenses. A fundamental solution to memory-based timing attacks is to eliminate the timing channel completely (e.g., by enforcing constant-time memory accesses or introducing intrusive and restrictive memory partitioning between different security contexts). However, eliminating the timing channel incurs very high performance overheads (as shown in Section 7). Therefore, several prior works [94, 95, 96, 97] propose defenses and detection mechanisms that restrict the use of timing channels with cache management-based methods, such as restricting the userspace applications’ access to cache management instructions (e.g., clflush instruction in x86 systems [94]) and detecting timing attacks based on cache access patterns (e.g., a high number of cache misses can indicate a timing attack) [95, 96, 97]. These defenses are inapplicable to PiM architectures because these architectures provide direct access to memory by bypassing the cache hierarchy.
@@
Based on our analysis, we introduce IMPACT, a set of high-throughput In-Memory Processing-based timing Attacks that leverage direct and fast main memory accesses enabled by PiM architectures. Using direct access to DRAM via PiM operations, IMPACT observes and exploits the shared DRAM row buffer states. We demonstrate that IMPACT achieves high communication and information leakage throughput by (i) eliminating expensive cache bypassing steps required by other main memory- and cache-based timing covert- and side-channel attacks, and (ii) leveraging the intrinsic parallelism of PiM operations. We showcase this finding on two covert- and one side-channel attacks. To protect future systems against IMPACT, we describe and analyze four defense mechanisms.
@@
IMPACT-PnM Covert Channel Variant. We build a covert channel exploiting a PnM architecture called PiM-enabled instructions [67] that enables executing instructions in compute units near DRAM. In this variant, the receiver uses PiM-enabled instructions to measure the time it takes to execute a simple arithmetic operation on a set of DRAM rows and detect row buffer conflicts. The sender transmits a message by encoding the message with row buffer conflicts in selected addresses and carefully creating conflicts with the receiver’s accesses using PiM-enabled instructions. By executing instructions near memory, IMPACT-PnM achieves 8.2 Mb/s communication throughput, which is 3.6$\times$ faster than the state-of-the-art main memory-based covert-channel attack [98].
IMPACT-PuM Covert Channel Variant. We build a covert channel exploiting a PuM technique called RowClone [75] that enables bulk data copy and initialization. In this variant, the sender transmits a message with in-memory row-copy operations by copying rows in different DRAM banks in parallel for different bits of the message. The receiver decodes the message by issuing row-copy operations and measuring their latency. By exploiting the parallelism in memory, IMPACT-PuM achieves a communication throughput of 14.8 Mb/s, which is 6.5$\times$ faster than the state-of-the-art [98].
@@
IMPACT Side-Channel Attack. We build a PnM-based IMPACT side-channel attack that exploits PiM-enabled instructions [67] to leak private information of a concurrently-running application. We showcase this side-channel attack on genomic read mapping, one of the fundamental applications in DNA sequence analysis where preserving the privacy of the human genomic data is crucial [34, 99, 36, 100, 101, 102, 103]. Our side-channel attack leverages PiM-enabled instructions [67] to leak the private characteristics of a user’s query genome by observing the memory access patterns of the read mapping application. Our evaluation demonstrates that our side-channel attack leaks the properties of a query genome at a throughput of 7.6 Mb/s with 96% accuracy.
@@
Defense Mechanisms. Based on our security analysis, we discuss and evaluate potential defense mechanisms to mitigate IMPACT attacks by eliminating the timing channel and reducing the attack throughput. We describe the performance versus security trade-offs of these defenses. We eventually conclude that mitigating IMPACT incurs high performance overheads and more research is needed to find low-overhead solutions.
@@
  • To our knowledge, this is the first work that analyzes and evaluates the security of emerging PiM architectures against timing covert-channel and side-channel attacks.
  • We introduce IMPACT, a new set of high-throughput main memory-based timing attacks that leverage PiM operations to gain fast and reliable access to main memory.
  • We demonstrate two IMPACT covert channel variants, IMPACT-PnM and IMPACT-PuM that exploit PnM and PuM techniques, respectively. IMPACT-PnM allows the sender and the receiver to exchange messages using PiM-enabled instructions. IMPACT-PuM further leverages the parallelism provided by bulk in-memory copy operations like RowClone. We show that both attacks outperform the state-of-the-art main-memory-based covert-channel attacks in terms of communication throughput.
  • We demonstrate an IMPACT side-channel attack on genomic read mapping by leveraging PnM, using PiM-enabled instructions. We show that our side-channel attack leaks sensitive information with high throughput and a low error rate.
  • We describe and evaluate the trade-offs of four defense mechanisms to protect PiM-enabled systems against IMPACT’s covert- and side-channel attacks. We call for future research to find low-overhead solutions.
@@ 2 Background
This section provides a concise overview of 1) DRAM organization and operation, 2) processing-in-memory, and 3) main memory-based timing attacks.
@@ 2.1 DRAM Organization and Operation
Organization. Fig. 1 shows the hierarchical organization of a modern DRAM-based main memory. A memory channel connects the processor (CPU) to a DRAM module, where a module consists of multiple DRAM ranks. A rank consists of multiple DRAM chips that operate in lockstep. Each DRAM chip contains multiple DRAM banks that can be accessed independently. A DRAM bank is organized as a two-dimensional array of DRAM cells, where a row of cells is called a DRAM row. A DRAM cell consists of 1) a storage capacitor, which stores one bit of information in the form of electrical charge, and 2) an access transistor, which connects the capacitor to the row buffer through a bitline controlled by a wordline.
@@
Operation. To perform a read or write operation, the memory controller first needs to open a row, i.e., copy the data of the cells in the row to the row buffer. To open a row, the memory controller issues an activation command to a bank by specifying the row address to open. After activation is complete, the memory controller issues either read or write commands to read or write a DRAM word within the activated row. Subsequent accesses to the same row can be served quickly without fetching the data to the row buffer, creating a row hit [104]. To access data from another DRAM row in the same bank, the memory controller must first close the currently open row by issuing a bank precharge command, creating a row conflict/miss.
@@ 2.2 Processing-in-Memory
Processing-in-Memory (PiM) moves computation closer to where the data resides, alleviating the data movement bottleneck between memory and the processor [13, 43, 8, 44, 45, 46, 11, 47, 19, 48, 1, 2]. There are two main PiM approaches [49, 10, 8]: 1) Processing-near-Memory (PnM) and 2) Processing-using-Memory (PuM). PnM places computation logic near memory arrays (e.g., in a DRAM chip, next to each bank, or at the logic layer of 3D-stacked memory [50, 51, 52, 53, 37, 54, 55, 38, 56, 57, 21, 58, 59, 60, 26]). The PnM logic can execute application regions (e.g., instructions [67, 53], functions [48], application threads [20, 77]) depending on the design. PnM computation units typically employ small caches due to 1) area constraints and 2) the ineffectiveness of caches in improving the performance of a wide range of memory-intensive applications. Therefore, various systems [53, 67, 48] offload the memory-intensive application regions on PnM units, while executing more cache-friendly application regions on the host CPU with a deeper cache hierarchy. PuM uses analog operational properties of the memory circuitry (e.g., DRAM) to enable massively parallel in-memory computation [70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85].
@@ 2.3 Main Memory Timing Attacks
Main memory timing attacks exploit several shared structures in the main memory. One such structure is the DRAM row buffer that acts as an internal cache inside DRAM banks. An attacker can understand whether a row is recently accessed (i.e., already in the row buffer) by measuring the time it takes to access that row. Using this observation, DRAMA [98] builds a side-channel attack to leak keystrokes in a real system. Several prior works (e.g., [98, 105, 106, 107, 108]) leak DRAM address mapping functions. Other main memory-based attacks exploit the memory bus contention [109] and memory deduplication [110, 111] and (de)compression [112, 113, 114, 115, 116, 117, 118, 119, 120] latencies.
@@ 3 A Case for PiM-based Timing Attacks
The adoption of PiM architectures has been gaining momentum in recent years because they provide high performance and low energy consumption by reducing data movement. Therefore, it is timely and critical to analyze and evaluate the security of PiM architectures to prevent any vulnerabilities from causing widespread susceptibility as PiM architectures become widely available. Unfortunately, no prior work analyzes and evaluates the security of emerging PiM architectures against timing covert- and side-channel attacks.
After a thorough analysis of existing PiM architectures, we show that adopting PiM solutions creates opportunities for high-throughput main memory-based timing attacks for two main reasons. First, the adoption of PiM solutions provides userspace applications with fast and reliable direct access to main memory to reduce data movement. Second, defenses against these attacks either incur high performance overheads or are not applicable to PiM architectures. In this work, we study the impact of PiM architectures on row buffer-based timing channels and their throughput.